# Nanocavities for Molecular Optomechanics: their fundamental description and applications

### v1

@@ Introduction
A subtlety that is probably unique to McOM is the existence of a quasi-continuum of dark plasmonic modes that may spectrally overlap with the bright mode of interest (for which all parameters above are computed). Sometime referred to as the ‘pseudomode’ in the regime of strong coupling 93, 94, this quasi-continuum is responsible for quenching the emission of a dipole placed in the near field; the factor $\beta$ is defined as the fraction of photons injected in the bright mode of interest vs. the total photon emission rate in the near-field that includes all quenching channels. 83 We explain how to compute $\beta$ (which is a priori different for Stokes and anti-Stokes sidebands) in the Appendix. We recall that the $\beta$-factor is a useful figure-of-merit for single-photon sources embedded in photonic structures such as waveguides and micropillars, where it quantifies the likelihood that a photon is emitted in the guided mode vs. the continuum of free-space modes and is intimately linked to the Purcell factor. 95
@@
A special class of systems to which the theory of McOM was employed are so-called picocavities, which are believed to be atomic-scale hotspots randomly created in laser-illuminated nanocavities during SERS measurements 38. Their phenomenology consists of a set of suddenly appearing new Raman peaks (the ‘picocavity signal’) under continuous excitation, whose fully correlated fluctuations point to a single-molecule origin. They coexist with the usual SERS peaks (the ‘nanocavity signal’) whose intensities generally remain weakly affected by the appearance of a picocavity. The observed linear increase of the anti-Stokes to Stokes intensity ratio of the picocavity peaks with laser power was consistent with the prediction of quantum backaction by the McOM theory 38. (As mentioned earlier, McOM provides a set of rate equations for quantum backaction that are fully equivalent to those for vibrational pumping in SERS. 60) This observation implies that the cooperativity $\mathcal{C}$ is comparable to the thermal occupancy of the vibrational mode $\bar{n}_{\text{th}}$.
@@
Under the assumptions used in Ref. 38 that the decay rates $\kappa$ and $\gamma$ and the intracavity photon number $n_{p}$ are identical for nanocavity and picocavity, and with Raman polarizability estimated from DFT calculations of the bare molecule, the observation of vibrational pumping is consistent with an effective picocavity mode volume $V_{\text{pico}}$ on the order of 1 nm3. Such values, which are two to three orders of magnitude smaller than those of nanocavities, have been also predicted by electromagnetic simulations 100, 101, 102 (Fig. 4). We stress here that McOM is not a theory intended to explain the formation of picocavities.
@@
If we now apply the scaling law presented above, which states that the cooperativity scales as $Q/V$ in McOM for filled cavities, we arrive at the intriguing prediction that the Raman signal from a picocavity could be two or three orders of magnitude stronger than the stable nanocavity signal. We emphasize here that this prediction already accounts for the fact that hundreds of molecules contribute to the nanocavity signal whereas just one does for the picocavity signal. Such giant Raman intensity enhancement has so far not been observed: experiments reported picocavity signals of similar magnitude as the nanocavity signal, which itself shows little modification during a picocavity event 38, 103, 104, 105, 106, 102, 107, 108. There could be multiple explanations for this apparent discrepancy: On the one hand, molecules in picocavities could possibly suffer from stronger quenching, reducing $\eta_{\text{rad}}$. On the other hand, a chemical enhancement mechanism could play a dominant role in explaining picocavity spectra, as suggested by DFT calculations that include just a few gold atoms around the molecule.107 Finally, it is believed that large field gradients at the scale of the molecule are an essential feature of picocavities 38, so that this type of coupling should be properly integrated in the McOM model, originally built around the dipole approximation. In any case, it would be a worthwhile undertaking to gain better understanding of all parameters that enter the expression of detected Raman photon flux in such systems (Fig. 3).
@@
  • The optical response of the nanocavity should be fully known in the frequency region covering the laser and Raman sideband. Experimentally, dark-field scattering spectroscopy provides valuable information at the single particle level, yet it is only sensitive to bright modes that efficiently scatter an incoming plane wave and does not suffice to reconstruct the near-field response that also depends on dark modes. It is also not able to distinguish the contributions from $\kappa_{\text{ex}}$ and $\kappa_{\text{0}}$ to the observed $Q$-factor. Numerical simulations can be helpful to obtain the decay rates $\kappa_{\text{ex}}$, $\kappa_{0}$, the mode volume $V$ and the bright mode branching ration $\beta$. Simulations are quite accurate for gaps larger than 1 nm but require a precise knowledge of the geometry and refractive index of the molecular spacer (for gap nanocavities). Usually, these parameters are fine-tuned in the simulations to reproduce the measured dark-field scattering spectra. 109 Electron beam methods offer nanoscale insights into the plasmonic modes, 110 but NPoM geometries and vertical nanogaps in general are not readily accessible.
@@
  • To obtain the coupling rate $g_{0}$ and single-photon cooperativity $\mathcal{C}_{0}$, knowledge of the Raman polarisability and the exact number of embedded molecules is required. In reality, each molecule has a different coupling rate depending on its exact location due to the inhomogeneity of the near-field, which can be treated numerically. 51 But the main source of uncertainty here is the single-molecule Raman polarisability, and more particularly, how much it is affected by ‘chemical’ enhancement for molecules bound to the metal. 111, 112, 113, 45 Accurate calculations of Raman cross-sections for large enough metal clusters coupled to small organic molecules remain difficult and they suggest that the chemical enhancement factor is usually large. 107 Alternatively, a more controllable thin material such as graphene 114 could be used as an intermediate intermediate layer between the plasmonic metal and the molecules, in an attempt to better separate chemical and electromagnetic enhancements.
@@
  • Last, but not least, the intracavity plasmon number $n_{p}$ must be estimated for a given pump power, which requires the additional knowledge of $\eta_{\text{in}}$. Focused cylindrical vectorial beams allow optimising the power coupled to specific nanocavity modes, 67, 86 but a quantitative knowledge of $\eta_{\text{in}}$ remains elusive. Obtaining this parameter from a simulation implies the use of a focused beam as background field, often much harder to implement than a plane wave excitation 115, in particular for nanocavities built on mirrors 116. Experimental measurements of $\eta_{\text{in}}$ or $n_{p}$ are challenging in the context of McOM and SERS. Interferometric scattering microscopy 117 was successfully employed to quantify the extinction of nanoantennas on glass 118, 119 and should therefore allow retrieving $\eta_{\text{in}}$ in the context of McOM, where most cavities are assembled on a mirror substrate, though. Alternatively, and as elaborated below, the experimental geometry can be modified to resemble more that of the simulated plane wave excitation.
@@
In addition to these challenges linked with the difficulties to fully characterise the near-field response of a gap nanocavity, another hurdle in the quest for McOM effects and applications is the limited tolerance of metallic nanocavities to large excitation powers. In Ref.51, for example, irreversible changes to the Raman spectra (which can be associated to optical ‘damage’ in this context, Fig. S23 in Ref. 51) represent a non-negligible portion of the studied power-dependent effect – even though much care was taken to limit this damage by short laser exposure time and optimized pulse duration. The most exciting predictions of McOM are realised in the regime where $\mathcal{C}=n_{p}\mathcal{C}_{0}>1$, so that the intracavity plasmon number $n_{p}$ must reach the largest possible values – at least be of order unity. To avoid excessive heating, this regime has been mostly explored under picosecond pulsed excitation. 46, 47, 48, 49, 51 This causes at least two types of difficulties. First, because $\eta_{\text{in}}$ is usually much smaller than one, most of the incoming laser power is not even coupling to the nanocavity mode of interest, but it can nevertheless be absorbed in the focal area and generate unwanted heat, hot-electrons, etc. Second, as $n_{p}$ increases inside the nanocavity, other nonlinearities may emerge and are hard to experimentally distinguish from potential optomechanical effects. In particular, both thermal effects 109, 120, 121 and electrons excited in the metal 122, 123, 124, 125 are expected to modify the transient plasmonic response on time scales down to picoseconds 126, 127, 128 – comparable to the vibrational decay rate that dictates the optomechanical response.
@@
In fact, it remains an open question whether optomechanical nonlinearities can be observed before the onset of plasmonic nonlinearities. In most studies of ultrafast plasmonics on metallic nanoparticles, measurable nonlinear response (such as resonance shift and broadening) is observed for pump fluences as low as $100~{}\mu$J/cm2 at off-resonant near-infrared wavelengths (see for example Refs. 126, 125, 128, among others). When using a high pulse repetition rate of $\sim 100$ MHz typical of Ti:Sapph oscillators, this fluence translates in an average power of $100~{}\mu$W/$\mu$m2. Considering a diffraction-limited spot-size in a confocal microscope with numerical aperture of 0.8-0.9, and that excitation is usually resonant with a plasmonic mode in McOM experiments, it means that few tens of microwatts – or even less – averaged power may be enough to drive a significant nonlinearity in the plasmonic response. In comparison, the highest average power used in Ref. 51 was 60 $\mu$W. The optomechanical effects being investigated should therefore be carefully deconvoluted from the nonlinear plasmonic response itself, motivating further work.
@@
  • In contrast with macroscopic and mesoscopic mechanical resonators, molecular vibrations feature markedly anharmonic potentials, in particular for modes that are localized to few molecular bonds 58. The standard McOM Hamiltonian, however, assumes a harmonic potential for the vibration. One recent theoretical work extends the formalism to account for vibrational anharmonicity in the case of a single, off-resonant molecule 129. When sufficiently sharp features are present in the LDOS of the nanocavity, this work predicts a regime of incoherent mechanical blockade, provided that only the Stokes transition to the first excited vibration can be efficiently pumped by SERS, while the second one is not. Such sharply changing LDOS can be achieved for example by the Fano lineshape of a hybrid plasmonic-dielectric cavity 130, 131. Another finding of Ref.129 is that anharmonicity has deep impacts on the regime of dynamical backaction amplification: it changes the threshold for coherent mechanical oscillations and lowers their amplitudes. A recent proposal also suggests that vibrational anharmonicity is a resource in single-molecule optomechanics for the production of antibunched photons through optomechanical blockade.132 It is likely, however, that such effects rapidly disappear in the experimentally relevant setting of many molecules, since the collective modes should become harmonic.132 Solving a model that includes anharmonicity and collective effects remains an open challenge. We also mention here the case of strongly driven picocavities, where vibrational pumping combined with anharmonicity can lead to qualitatively similar Raman peak shifts as predicted from the optical spring calculation. 51
  • Primo et al. put forward that dissipative optomechanical coupling can be on par with dispersive coupling in McOM.90 Dissipative coupling occurs when the mechanical displacement modifies one of the dissipation rates of the cavity ($\kappa_{\text{ex}}$ and/or $\kappa_{0}$) 133, 99. The formula for $g_{0}$ originally derived in Refs. 5, 7 does not capture dissipative coupling, which is intimately related with the multimode and non-hermitian nature of nanocavities.134 The presence of dissipative coupling impacts the shape and magnitude of the optomechanical damping rate vs. laser wavelength 133, 99. Ref. 90 applies perturbation theory in the quasi-normal mode formalism to compute an imaginary part of the optomechanical coupling rate that corresponds to dissipative coupling, confirming the results with full-wave moving-mesh simulations. They conclude that dissipative coupling is at least of comparable magnitude as dispersive coupling in the canonical NPoM nanocavity, which we confirm using a simplified approach in Fig. 4b. It is also expected to prominently feature in engineered multimode McOM systems, like hybrid nanoantenna-cavity resonators.73
@@
  • Resonant and near-resonant Raman scattering could in first approximation be thought of as just increasing the Raman polarizability that enters the McOM model desribed in the first part. But recent theoretical work by M. Martinez-Garcia et al. concluded that, even in the weak optomechanical coupling regime, interference between resonant and nonresonant contributions to SERS can deeply alter the predicted nanocavity scattering spectrum.135 Previously, S. Hughes et al. already investigated the quantum nonlinear regime of resonant strong coupling between the nanocavity mode and the molecular electronic transition, in the presence of vibrational modes 136. Future experimental and theoretical works are needed clarify the precise impact of molecular electronic resonance on vibrational dynamics and nanocavity scattering spectrum in the context of accessible SERS scenarios.
  • As already discussed above, experimentally determining the intracavity plasmon number remains an open challenge; conversely, numerical estimates are usually performed under the simplifying assumption of plane wave excitation and there is little quantitative work that accounts for strongly focused laser beams, possibly with complex vectorial fields 86. Such simulations do exist 115, 116 but have not been implemented yet in the context of McOM to obtain more accurate estimates of input coupling efficiency and intracavity plasmon number.
  • A resource-efficient formalism that correctly accounts for collective effects among molecules coupled to a same nanocavity is still missing. The pioneering calculations presented in Ref. 51 require numerical evaluation using a discrete set of point dipoles representing the molecules, limiting this approach to 100 molecules so far, while typical nanocavities can host 1000 or more. Since 2D materials can also be used as Raman active layers in McOM 48, 49, the development of a continuum phonon model that applies also for molecular layers and accounts for collective dynamics would be valuable. Additionally, it was recently proposed that dipole-dipole coupling among closely packed IR-active molecules is an important source of collective behavior 137, 138 having impacts on the Raman spectra 139; it would be appealing to include such effects in a comprehensive McOM description.
@@
Figure 5: Optomechanically induced transparency: concept and measurement schemes. (a) Frequency representation of all involved fields and resonances, with the definitions of detunings $\Delta$ and $\delta$. (b) Same parameters represented on an energy level diagram in the single excitation sub-space. The notation $|p,\nu\rangle$ designates a state with $p$ plasmons and $\nu$ vibrations. (c) Computed map of OMIT signal strength (normalised to the input probe power) vs. cooperativity $\mathcal{C}$ and external coupling efficiency $\eta_{\text{rad}}$. (d) Possible OMIT measurement scheme in macroscopic cOM, adapted from Ref. 140. PLL: phase-locked loop; FC: filter cavity; PM: phase modulator. (e) Proposed implementation of OMIT measurement in McOM using a dark-field geometry for pump and probe excitation. A spectral filter blocks the pump and the probe signal is sent to the detector. The pump (pulsed or cw) is intensity-modulated so that the demodulated scattered probe intensity senses the pump-induced change in cavity scattering, which can have thermal (broad plasmonic response) or optomechanical (narrow vibrational response) origins.
@@
Contrary to macroscopic cavity optomechanics, where heterodyne linear optical field measurements are the norm, McOM has been so far only studied with photon-counting detectors such as CCD cameras, possibly losing valuable phase information in the process. In particular, depending on the detuning between laser and cavity resonance the optomechanical (Raman) sidebands are predicted to consist of different contents of phase and amplitude modulations: Pure phase modulation is expected on resonance, which can be readout by homodyne or heterodyne detection. We therefore suggest to introduce the use of local oscillators generated by a nonlinear optical process in the measurement of SERS signals from nanocavities, as already implemented in the case of stimulated Raman scattering (SRS) imaging 141 and vibrational sum frequency generation 36.
@@
Related, a valuable measurement that was not yet implemented as such in McOM is that of optomechanically induced transparency, or OMIT 142, 143. Beyond its fundamental appeal and potential applications 144, OMIT measurements allow a very direct estimate of the cooperativity, a capability that is still lacking in McOM. The effect is described in Fig. 5a,b: a strong pump laser is tuned on the lower vibrational sideband of the cavity, so that the anti-Stokes frequency is near resonant with the cavity frequency (their detuning is denoted as $\Delta$). A weaker probe laser beam is swept in frequency (detuning $\delta$ from the cavity). In macroscopic cOM, a destructive interference between the two excitation pathways depicted in Fig. 5b can induce a transparency window for the probe when $\Delta\simeq\delta$ compared to its absorption by the cavity in the absence of pump field. A detailed calculation adapted to the specificity of McOM (where scattered light is detected instead of reflection or transmission) has not been published to our knowledge; however we expect that a sharp feature is expected in the probe scattering spectrum (peak or dip) at $\Delta\simeq\delta$. Panel (c) in Fig. 5 estimates the relative strength of this feature (for the case $\Delta=0$) as a function of cooperativity $\mathcal{C}={n}_{p}\mathcal{C}_{0}$ (where ${n}_{p}$ scales linearly with pump power) and external coupling efficiency $\kappa_{\text{ex}}/\kappa=\eta_{\text{rad}}$. Two regimes are favorable to the observation of OMIT: (i) overcoupled cavity ($\eta_{\text{rad}}\sim 1$) with moderate cooperativity ($\mathcal{C}\simeq 1$), or more conventionally (ii) high cooperativity ($\mathcal{C}\gg 1$) under critical coupling ($\eta_{\text{rad}}=0.5$).
@@
A typical OMIT measurement scheme in macroscopic cOM is presented in Fig. 5d; how should OMIT be measured in the context of McOM on a plasmonic nanocavity? The first challenge is that plasmonic nanocavities are not probed in reflection or transmission given their subwavelength dimensions, but rather in scattering geometry. What matters is to isolate the optical field that has interacted with the nanocavity mode. For that, we propose a dark-field scattering detection geometry (Fig. 5e) where the specular reflection of pump and probe beam are not collected. The pump frequency is filtered out of scattered field and the probe power is monitored; modulating the pump intensity and demodulating the probe signal using lock-in detection should allow to reach the required sensitivity for observing OMIT with state-of-the-art nanocavities (Fig. 5c). Compared to recent implementations of SRS on plasmonic antennas 145, 146, the proposed scheme mainly differs in the dark-field detection geometry and the prerequisite to be able to observe the absorption or scattering of a single nanocavity; also, note that the pump and probe beams in OMIT correspond to the Stokes and pump beams in SRS, respectively. This measurement could be performed under ps pulsed or continuous wave excitation.

### v2

@@ Introduction
A subtlety that is probably unique to McOM is the existence of a quasi-continuum of dark plasmonic modes that may spectrally overlap with the bright mode of interest (for which all parameters above are computed). Sometime referred to as the ‘pseudomode’ in the regime of strong coupling 93, 94, this quasi-continuum is responsible for quenching the emission of a dipole placed in the near field; the $\beta$-factor is defined as the fraction of photons injected in the bright mode of interest vs. the total photon emission rate in the near-field that includes all quenching channels. 83 We explain how to compute $\beta$ (which is a priori different for Stokes and anti-Stokes sidebands) in the Appendix. We recall that the $\beta$-factor is a useful figure-of-merit for single-photon sources embedded in photonic structures such as waveguides and micropillars, where it quantifies the likelihood that a photon is emitted in the guided mode vs. the continuum of free-space modes and is intimately linked to the Purcell factor. 95
@@
A special class of systems to which the theory of McOM was employed are so-called picocavities, which are believed to be atomic-scale hotspots randomly created in laser-illuminated nanocavities during SERS measurements 38. Their phenomenology consists of a set of suddenly appearing new Raman peaks (the ‘picocavity signal’) under continuous excitation, whose fully correlated fluctuations point to a single-molecule origin. They coexist with the usual SERS peaks (the ‘nanocavity signal’) whose intensities generally remain weakly affected by the appearance of a picocavity. The observed linear increase of the anti-Stokes to Stokes intensity ratio of the picocavity peaks with laser power was consistent with the prediction of quantum backaction by the McOM theory 38. (As mentioned earlier, McOM provides a set of rate equations for quantum backaction that are fully equivalent to those for vibrational pumping in SERS. 60) This observation implies that the cooperativity $\mathcal{C}$ is comparable to the thermal occupancy of the vibrational mode ${n}_{\text{th}}$.
@@
Under the assumptions used in Ref. 38 that the decay rates $\kappa$ and $\gamma$ and the intracavity photon number $n_{p}$ are identical for nanocavity and picocavity, and with Raman polarizability estimated from DFT calculations of the bare molecule, the observation of vibrational pumping is consistent with an effective picocavity mode volume $V_{\text{pico}}$ on the order of 1 nm3. Such values, which are two to three orders of magnitude smaller than those of nanocavities, have been also predicted by electromagnetic simulations 100, 101, 102, 103 (Fig. 4). We stress here that McOM is not a theory intended to explain the formation of picocavities.
@@
If we now apply the scaling law presented above, which states that the cooperativity scales as $Q/V$ in McOM for filled cavities, we arrive at the intriguing prediction that the Raman signal from a picocavity could be two or three orders of magnitude stronger than the stable nanocavity signal. We emphasize here that this prediction already accounts for the fact that hundreds of molecules contribute to the nanocavity signal whereas just one does for the picocavity signal. Such giant Raman intensity enhancement has so far not been observed: experiments reported picocavity signals of similar magnitude as the nanocavity signal, which itself shows little modification during a picocavity event 38, 104, 105, 106, 107, 103, 108, 109. There could be multiple explanations for this apparent discrepancy: On the one hand, molecules in picocavities could possibly suffer from stronger quenching, reducing $\eta_{\text{rad}}$. On the other hand, a chemical enhancement mechanism could play a dominant role in explaining picocavity spectra, as suggested by DFT calculations that include just a few gold atoms around the molecule.108 Finally, it is believed that large field gradients at the scale of the molecule are an essential feature of picocavities 38, so that this type of coupling should be properly integrated in the McOM model, originally built around the dipole approximation. In any case, it would be a worthwhile undertaking to gain better understanding of all parameters that enter the expression of detected Raman photon flux in such systems (Fig. 3).
@@
  • The optical response of the nanocavity should be fully known in the frequency region covering the laser and Raman sideband. Experimentally, dark-field scattering spectroscopy provides valuable information at the single particle level, yet it is only sensitive to bright modes that efficiently scatter an incoming plane wave and does not suffice to reconstruct the near-field response that also depends on dark modes. It is also not able to distinguish the contributions from $\kappa_{\text{ex}}$ and $\kappa_{\text{0}}$ to the observed $Q$-factor. Numerical simulations can be helpful to obtain the decay rates $\kappa_{\text{ex}}$, $\kappa_{0}$, the mode volume $V$ and the bright mode branching ratio $\beta$. Simulations are quite accurate for gaps larger than 1 nm but require a precise knowledge of the geometry and refractive index of the molecular spacer (for gap nanocavities). Usually, these parameters are fine-tuned in the simulations to reproduce the measured dark-field scattering spectra. 110 Electron beam methods offer nanoscale insights into the plasmonic modes, 111 but NPoM geometries and vertical nanogaps in general are not readily accessible.
@@
  • To obtain the coupling rate $g_{0}$ and single-photon cooperativity $\mathcal{C}_{0}$, knowledge of the Raman polarisability and the exact number of embedded molecules is required. In reality, each molecule has a different coupling rate depending on its exact location due to the inhomogeneity of the near-field, which can be treated numerically. 51 But the main source of uncertainty here is the single-molecule Raman polarisability, and more particularly, how much it is affected by ‘chemical’ enhancement for molecules bound to the metal. 112, 113, 114, 45 Accurate calculations of Raman cross-sections for large enough metal clusters coupled to small organic molecules remain difficult and they suggest that the chemical enhancement factor is usually large. 108 Alternatively, a more controllable thin material such as graphene 115 could be used as an intermediate intermediate layer between the plasmonic metal and the molecules, in an attempt to better separate chemical and electromagnetic enhancements.
@@
  • Last, but not least, the intracavity plasmon number $n_{p}$ must be estimated for a given pump power, which requires the additional knowledge of $\eta_{\text{in}}$. Focused cylindrical vectorial beams allow optimising the power coupled to specific nanocavity modes, 67, 86 but a quantitative knowledge of $\eta_{\text{in}}$ remains elusive. Obtaining this parameter from a simulation implies the use of a focused beam as background field, often much harder to implement than a plane wave excitation 116, in particular for nanocavities built on mirrors 117. Experimental measurements of $\eta_{\text{in}}$ or $n_{p}$ are challenging in the context of McOM and SERS. Interferometric scattering microscopy 118 was successfully employed to quantify the extinction of nanoantennas on glass 119, 120 and should therefore allow retrieving $\eta_{\text{in}}$ in the context of McOM, where most cavities are assembled on a mirror substrate, though. Alternatively, and as elaborated below, the experimental geometry can be modified to resemble more that of the simulated plane wave excitation.
@@
In addition to these challenges linked with the difficulties to fully characterise the near-field response of a gap nanocavity, another hurdle in the quest for McOM effects and applications is the limited tolerance of metallic nanocavities to large excitation powers. In Ref.51, for example, irreversible changes to the Raman spectra (which can be associated to optical ‘damage’ in this context, Fig. S23 in Ref. 51) represent a non-negligible portion of the studied power-dependent effect – even though much care was taken to limit this damage by short laser exposure time and optimized pulse duration. The most exciting predictions of McOM are realised in the regime where $\mathcal{C}=n_{p}\mathcal{C}_{0}>1$, so that the intracavity plasmon number $n_{p}$ must reach the largest possible values – at least be of order unity. To avoid excessive heating, this regime has been mostly explored under picosecond pulsed excitation. 46, 47, 48, 49, 51 This causes at least two types of difficulties. First, because $\eta_{\text{in}}$ is usually much smaller than one, most of the incoming laser power is not even coupling to the nanocavity mode of interest, but it can nevertheless be absorbed in the focal area and generate unwanted heat, hot-electrons, etc. Second, as $n_{p}$ increases inside the nanocavity, other nonlinearities may emerge and are hard to experimentally distinguish from potential optomechanical effects. In particular, both thermal effects 110, 121, 122 and electrons excited in the metal 123, 124, 125, 126 are expected to modify the transient plasmonic response on time scales down to picoseconds 127, 128, 129 – comparable to the vibrational decay rate that dictates the optomechanical response.
@@
In fact, it remains an open question whether optomechanical nonlinearities can be observed before the onset of plasmonic nonlinearities. In most studies of ultrafast plasmonics on metallic nanoparticles, measurable nonlinear response (such as resonance shift and broadening) is observed for pump fluences as low as $100~{}\mu$J/cm2 at off-resonant near-infrared wavelengths (see for example Refs. 127, 126, 129, among others). When using a high pulse repetition rate of $\sim 100$ MHz typical of Ti:Sapph oscillators, this fluence translates in an average power of $100~{}\mu$W/$\mu$m2. Considering a diffraction-limited spot-size in a confocal microscope with numerical aperture of 0.8-0.9, and that excitation is usually resonant with a plasmonic mode in McOM experiments, it means that few tens of microwatts – or even less – averaged power may be enough to drive a significant nonlinearity in the plasmonic response. In comparison, the highest average power used in Ref. 51 was 60 $\mu$W. The optomechanical effects being investigated should therefore be carefully deconvoluted from the nonlinear plasmonic response itself, motivating further work.
@@
  • In contrast with macroscopic and mesoscopic mechanical resonators, molecular vibrations feature markedly anharmonic potentials, in particular for modes that are localized to few molecular bonds 58. The standard McOM Hamiltonian, however, assumes a harmonic potential for the vibration. One recent theoretical work extends the formalism to account for vibrational anharmonicity in the case of a single, off-resonant molecule 130. When sufficiently sharp features are present in the LDOS of the nanocavity, this work predicts a regime of incoherent mechanical blockade, provided that only the Stokes transition to the first excited vibration can be efficiently pumped by SERS, while the second one is not. Such sharply changing LDOS can be achieved for example by the Fano lineshape of a hybrid plasmonic-dielectric cavity 131, 132. Another finding of Ref.130 is that anharmonicity has deep impacts on the regime of dynamical backaction amplification: it changes the threshold for coherent mechanical oscillations and lowers their amplitudes. A recent proposal also suggests that vibrational anharmonicity is a resource in single-molecule optomechanics for the production of antibunched photons through optomechanical blockade.133 It is likely, however, that such effects rapidly disappear in the experimentally relevant setting of many molecules, since the collective modes should become harmonic.133 Solving a model that includes anharmonicity and collective effects remains an open challenge. We also mention here the case of strongly driven picocavities, where vibrational pumping combined with anharmonicity can lead to qualitatively similar Raman peak shifts as predicted from the optical spring calculation. 51
  • Primo et al. put forward that dissipative optomechanical coupling can be on par with dispersive coupling in McOM.90 Dissipative coupling occurs when the mechanical displacement modifies one of the dissipation rates of the cavity ($\kappa_{\text{ex}}$ and/or $\kappa_{0}$) 134, 99. The formula for $g_{0}$ originally derived in Refs. 5, 7 does not capture dissipative coupling, which is intimately related with the multimode and non-hermitian nature of nanocavities.135 The presence of dissipative coupling impacts the shape and magnitude of the optomechanical damping rate vs. laser wavelength 134, 99. Ref. 90 applies perturbation theory in the quasi-normal mode formalism to compute an imaginary part of the optomechanical coupling rate that corresponds to dissipative coupling, confirming the results with full-wave moving-mesh simulations. They conclude that dissipative coupling is at least of comparable magnitude as dispersive coupling in the canonical NPoM nanocavity, which we confirm using a simplified approach in Fig. 4b. It is also expected to prominently feature in engineered multimode McOM systems, like hybrid nanoantenna-cavity resonators.73
@@
  • Resonant and near-resonant Raman scattering could in first approximation be thought of as just increasing the Raman polarizability that enters the McOM model desribed in the first part. But recent theoretical work by M. Martinez-Garcia et al. concluded that, even in the weak optomechanical coupling regime, interference between resonant and nonresonant contributions to SERS can deeply alter the predicted nanocavity scattering spectrum.136 Previously, S. Hughes et al. already investigated the quantum nonlinear regime of resonant strong coupling between the nanocavity mode and the molecular electronic transition, in the presence of vibrational modes 137. Future experimental and theoretical works are needed clarify the precise impact of molecular electronic resonance on vibrational dynamics and nanocavity scattering spectrum in the context of accessible SERS scenarios.
  • As already discussed above, experimentally determining the intracavity plasmon number remains an open challenge; conversely, numerical estimates are usually performed under the simplifying assumption of plane wave excitation and there is little quantitative work that accounts for strongly focused laser beams, possibly with complex vectorial fields 86. Such simulations do exist 116, 117 but have not been implemented yet in the context of McOM to obtain more accurate estimates of input coupling efficiency and intracavity plasmon number.
  • A resource-efficient formalism that correctly accounts for collective effects among molecules coupled to a same nanocavity is still missing. The pioneering calculations presented in Ref. 51 require numerical evaluation using a discrete set of point dipoles representing the molecules, limiting this approach to 100 molecules so far, while typical nanocavities can host 1000 or more. Since 2D materials can also be used as Raman active layers in McOM 48, 49, the development of a continuum phonon model that applies also for molecular layers and accounts for collective dynamics would be valuable. Additionally, it was recently proposed that dipole-dipole coupling among closely packed IR-active molecules is an important source of collective behavior 138, 139 having impacts on the Raman spectra 140; it would be appealing to include such effects in a comprehensive McOM description.
@@
Figure 5: Optomechanically induced transparency: concept and measurement schemes. (a) Frequency representation of all involved fields and resonances, with the definitions of detunings $\Delta$ and $\delta$. (b) Same parameters represented on an energy level diagram in the single excitation sub-space. The notation $|p,\nu\rangle$ designates a state with $p$ plasmons and $\nu$ vibrations. (c) Computed map of OMIT signal strength (normalised to the input probe power) vs. cooperativity $\mathcal{C}$ and external coupling efficiency $\eta_{\text{rad}}$. (d) Possible OMIT measurement scheme in macroscopic cOM, adapted from Ref. 141. PLL: phase-locked loop; FC: filter cavity; PM: phase modulator. (e) Proposed implementation of OMIT measurement in McOM using a dark-field geometry for pump and probe excitation. A spectral filter blocks the pump and the probe signal is sent to the detector. The pump (pulsed or cw) is intensity-modulated so that the demodulated scattered probe intensity senses the pump-induced change in cavity scattering, which can have thermal (broad plasmonic response) or optomechanical (narrow vibrational response) origins.
@@
Contrary to macroscopic cavity optomechanics, where heterodyne linear optical field measurements are the norm, McOM has been so far only studied with photon-counting detectors such as CCD cameras, possibly losing valuable phase information in the process. In particular, depending on the detuning between laser and cavity resonance the optomechanical (Raman) sidebands are predicted to consist of different contents of phase and amplitude modulations: Pure phase modulation is expected on resonance, which can be readout by homodyne or heterodyne detection. We therefore suggest to introduce the use of local oscillators generated by a nonlinear optical process in the measurement of SERS signals from nanocavities, as already implemented in the case of stimulated Raman scattering (SRS) imaging 142 and vibrational sum frequency generation 36.
@@
Related, a valuable measurement that was not yet implemented as such in McOM is that of optomechanically induced transparency, or OMIT 143, 144. Beyond its fundamental appeal and potential applications 145, OMIT measurements allow a very direct estimate of the cooperativity, a capability that is still lacking in McOM. The effect is described in Fig. 5a,b: a strong pump laser is tuned on the lower vibrational sideband of the cavity, so that the anti-Stokes frequency is near resonant with the cavity frequency (their detuning is denoted as $\Delta$). A weaker probe laser beam is swept in frequency (detuning $\delta$ from the cavity). In macroscopic cOM, a destructive interference between the two excitation pathways depicted in Fig. 5b can induce a transparency window for the probe when $\Delta\simeq\delta$ compared to its absorption by the cavity in the absence of pump field. A detailed calculation adapted to the specificity of McOM (where scattered light is detected instead of reflection or transmission) has not been published to our knowledge; however we expect that a sharp feature is expected in the probe scattering spectrum (peak or dip) at $\Delta\simeq\delta$. Panel (c) in Fig. 5 estimates the relative strength of this feature (for the case $\Delta=0$) as a function of cooperativity $\mathcal{C}={n}_{p}\mathcal{C}_{0}$ (where ${n}_{p}$ scales linearly with pump power) and external coupling efficiency $\kappa_{\text{ex}}/\kappa=\eta_{\text{rad}}$. Two regimes are favorable to the observation of OMIT: (i) overcoupled cavity ($\eta_{\text{rad}}\sim 1$) with moderate cooperativity ($\mathcal{C}\simeq 1$), or more conventionally (ii) high cooperativity ($\mathcal{C}\gg 1$) under critical coupling ($\eta_{\text{rad}}=0.5$).
@@
A typical OMIT measurement scheme in macroscopic cOM is presented in Fig. 5d; how should OMIT be measured in the context of McOM on a plasmonic nanocavity? The first challenge is that plasmonic nanocavities are not probed in reflection or transmission given their subwavelength dimensions, but rather in scattering geometry. What matters is to isolate the optical field that has interacted with the nanocavity mode. For that, we propose a dark-field scattering detection geometry (Fig. 5e) where the specular reflection of pump and probe beam are not collected. The pump frequency is filtered out of scattered field and the probe power is monitored; modulating the pump intensity and demodulating the probe signal using lock-in detection should allow to reach the required sensitivity for observing OMIT with state-of-the-art nanocavities (Fig. 5c). Compared to recent implementations of SRS on plasmonic antennas 146, 147, the proposed scheme mainly differs in the dark-field detection geometry and the prerequisite to be able to observe the absorption or scattering of a single nanocavity; also, note that the pump and probe beams in OMIT correspond to the Stokes and pump beams in SRS, respectively. This measurement could be performed under ps pulsed or continuous wave excitation.